OpenStudy (anonymous):

find an expression for the function whose graph is the given curve. for the line segment joining the points (-5, 10) and (7, -10)

OpenStudy (anonymous):

for the line segment joining the points (-5, 10) and (7, -10)

OpenStudy (ghazi):

(y-10)= (-10-10)/(7+5)*(x+5)

OpenStudy (anonymous):

so the expression will be y=(-5/3)x + (50/3)

OpenStudy (anonymous):

how do i account for only the segment?

OpenStudy (ghazi):

(y-10)=(-20)/(12)*(x+5)

OpenStudy (ghazi):

\[(y-10)=\frac{ -20 }{ 12 }*(x+5)\]